OpenStudy (anonymous):

The graph of y = ax2 + bx + c is a parabola that opens up and has a vertex at (-2, 5). What is the solution set of the related equation 0 = ax2 + bx + c?

OpenStudy (tad1):

-2 = -b/2a. Therefore b = 4a. Substitute into 5 = 4a -2b +c and it becomes -b + c = 5. We also have the equation 0 = 2ax + bx + c = x(2a + b) + c = x((3/2b)x + c So we have a system of two equations in two unknowns which can be solved and then back-substituted to find a. I'm sure there's a cleaner way to do this problem, but this way works.